hartnn (hartnn):

@banumati234 Prove that the vertically opposite angles of 2 intersecting lines are equal.

OpenStudy (anonymous):

Using what assumptions?

hartnn (hartnn):

this was @banumati234 's question, she could not post it. let here come here

hartnn (hartnn):

ok, you know what are supplementary angles ?

OpenStudy (anonymous):

yes.when the sum of 2 angles =180 degree,then the angles are supplementary

hartnn (hartnn):

good, so,  i marked 4 angles, can you tell me which angles are supplemenatry ?

OpenStudy (anonymous):

angle A,angle B angle C,angle D angle D,angle A angle B,angle C

OpenStudy (anonymous):

to prove that angle A=angle C

hartnn (hartnn):

oh, excellent! so, let angle A = x what can you say about angle B ?

OpenStudy (anonymous):

180-x

hartnn (hartnn):

to prove : A=C and B=D

OpenStudy (anonymous):

yes..

hartnn (hartnn):

right, since B and C are also supplementary, and B= 180-x what can you say about angle C ?

OpenStudy (anonymous):

angle C=x ok so angle A=angle C=x thanks!!!!!

hartnn (hartnn):

correct! also, since C and D are supplementary, D =180-x which = B so, B=D

OpenStudy (anonymous):

thanks again...

hartnn (hartnn):

welcome ^_^